OpenStudy (anonymous):

find the equation of the axis of symmetry of y=2x^2+6x+7

hero (hero):

axis of symmetry: x = -b/2a

hero (hero):

a = 2 b = 6 x =-6/2(2) x = -6/4 x =-2/3

OpenStudy (anonymous):

thanks! so the answer is: x= -2/3?

hero (hero):

Yes

hero (hero):

You can thank me with a medal

OpenStudy (anonymous):

okay.. thanks you!